12.( ) What is the future value of a $10,000 college tuition fund if the nominal rate of interest is 12 percent compounded quarterly for five years? a. $17,623. 42 b. $18,166. 97 c. $16,105.10 d. $18,061.11
Answer #1

Answer is d) $18,061.11

Explanation:

TO find future vale we use the below formula

A=P(1+r/n)Power(n*t)

Where, A = Future Value
P = principal or amount of money deposited,
r = annual interest rate
n = number of times compounded per year
t = time in years

In the given question we have,

P=$10000 , r=12% , n=4 and t=5 years

After plugging the given information we have

Future Value =$10000(1+0.12/4)Power 4*5=10000*1.806111=$18061.11